the sum if four times a number and three is the same as the difference if twice the number and eleven. Find the number.
: the sum if four times a number and three is the same as the difference if twice the number and eleven. Find the number.

Answer by checkley77(1452)

4X+3=2X-11
4X-2X=-11-3
2X=-14
X=-14/2
X=-7 ANSWER.
PROOF:
4*-7+3=2*-7-11
-28+3=-14-11
-25=-25